Prove that if n is an odd positive integer, then n^2 ≡ 1 (mod 8).
1
Expert's answer
2020-10-20T13:19:23-0400

n is an odd positive integer. We take n=2k+1, where k"\geq 0"


"n^{2}-1=(n+1)(n-1)=" (2k+2)2k=4(k+1)k

Note that k(k+1) is always an even number for any value of k.

Therefore k(k+1)=2 m

This imply 8"|n^{2}-1"

Hence "n^{2}\cong 1(mod 8)"